1) Find an equation of the line.Write the equation in standard form.
With slope of -4/7, through (5,3)
a) 7x+4y = -41
b) c) 4x + 7y = 41
c) 4x + 7y = -41
2) Find the slope of the line.
through (-3,3) and (-8,-9)
a. 6/11
b. 12/5
c.5/12
3) Find the slope of the line.
7x +y = -1
a. -1/7
b. 7
c. -7
4) Determine the slope and the y-intercept of the graph of 10x - 8y = 80
a. 4/5; (0,8)
b. -5/4; (0,-10)
c. 5/4; (0,-10)
5) Find an equation for the line.Write the equation in standard form.
Through the origin and (-8,-5)
a. -5x - 8y = 0
b. 5x - 8y = 0
c. -8x + 5y = 0



Answer by stanbon(75887)

1) Find an equation of the line.Write the equation in standard form.
With slope of -4/7, through (5,3)
Form: y = mx+b
3 = (-4/7)5 + b
21/7 + 28/7 = b
b = 7
------
Equation:
y = (-4/7)x + 7
4x + 7y = 49
a) 7x+4y = -41
b) c) 4x + 7y = 49
c) 4x + 7y = -41
---------------------------
2) Find the slope of the line.
through (-3,3) and (-8,-9)
slope = (3--9)/(-3--8) = 12/5
a. 6/11
b. 12/5
c.5/12
---------------------------------
3) Find the slope of the line.
7x +y = -1
y = -7x - 1
slope = -7
a. -1/7
b. 7
c. -7
4) Determine the slope and the y-intercept of the graph of 10x - 8y = 80
a. 4/5; (0,8)
b. -5/4; (0,-10)
c. 5/4; (0,-10)
5) Find an equation for the line.
Write the equation in standard form.
Through the origin and (-8,-5)
y = (5/8)x
5x - 8y = 0
a. -5x - 8y = 0
b. 5x - 8y = 0
c. -8x + 5y = 0
======================
Cheers,
Stan H.
======================